OpenStudy (anonymous):

Given: 1/m-4 = 2/m-2 a. What values should be excluded? b. Solve the equation.

OpenStudy (anonymous):

@prowrestler

OpenStudy (anonymous):

first cross multiply 1(m-2)=2(m-4) expand the bracket m-2=2m-8 collect like terms m-2m=-8+2 simplify -m=-6 divide both side by -1 m=6

OpenStudy (anonymous):

@Mariam.A What values should be excluded?